# Comparative Performance Evaluation of 5G-TSN Applications in Indoor Factory Environments

### v1

@@ V Conclusion
The move towards Industry 4.0 necessitates that factories replace their current industrial networks, which rely on Ethernet and other wired technologies, with wireless networks. This transition is essential for achieving the desired flexibility and adaptability in manufacturing environments. While some major companies have begun to experiment with 5G wireless communication technology, deployments are still at an early stage of development. In this study, we have simulated a 5G-TSN network deployed in an indoor factory to assess its suitability for industrial applications. We conducted simulations of 5G-TSN across various factory environments to understand its behaviour under different wireless channel profiles. Our results indicate that 5G-TSN can effectively handle latency-critical applications in indoor factories in a controlled environment, suggesting its potential for significant performance improvements. Since no prior work has been done in this specific context, our research serves as a baseline for evaluating 5G-TSN in factory scenarios. However, we plan to extend this work by testing more realistic scenarios with real-world parameters, such as varying numbers of UEs, gNB, and other dynamic factors including heavy network load, and more dynamic environmental interference. These tests will allow us to assess the scalability, reliability, and adaptability of the proposed approach in more complex and dynamic indoor factory environments in a more comprehensive manner. Based on the findings of this study, future work could also focus on real-world implementation and testing of 5G-TSN in industrial environments to validate the simulation results and identify any challenges or limitations. Additionally, research could explore the integration of 5G-TSN with existing factory automation systems and infrastructure to ensure seamless compatibility and maximize its benefits.

### v2

@@ V Conclusion
The move towards Industry 4.0 necessitates that factories replace their current industrial networks, which rely on Ethernet and other wired technologies, with wireless networks. This transition is essential for achieving the desired flexibility and adaptability in manufacturing environments. While some companies have begun to experiment with 5G technology, deployments are still at an early stage of development. In this study, we have simulated a 5G-TSN network deployed in an indoor factory to assess its suitability for industrial applications. We conducted simulations of 5G-TSN across various factory environments to understand its behaviour under different wireless channel profiles. Our results indicate that 5G-TSN can effectively handle latency-critical applications in indoor factories in a controlled environment, suggesting its potential for significant performance improvements. Since no prior work has been done in this specific context, our research serves as a baseline for evaluating 5G-TSN in factory scenarios. However, we plan to extend this work by testing more realistic scenarios with real-world parameters, such as varying numbers of UEs, gNB, varying network load, and more dynamic environmental interference. These tests will allow us to assess the scalability, reliability, and adaptability of the proposed approach in more complex and dynamic indoor factory environments in a more comprehensive manner. Based on the findings of this study, future work could also focus on real-world implementation and testing of 5G-TSN in industrial environments to validate the simulation results and identify any challenges or limitations. Additionally, research could explore the integration of 5G-TSN with existing factory automation systems and infrastructure to ensure seamless compatibility and maximize its benefits.
Acknowledgments
This publication has emanated from research conducted with the financial support of Research Ireland under Grant number 13/RC/2077_P2.